OpenStudy (anonymous):

I don't understand what Problem 3 pset3 is asking me to do. Would someone mind explaining?

OpenStudy (anonymous):

I'm on the pset about the biological perspective.

OpenStudy (anonymous):

Could you be a bit more specific? Did you read the assignment entirely?

OpenStudy (anonymous):

n = (1, 5, 8, 15, 29) k = (3, 9, 13, 19, 36) m = 4 ? n + m + 1 = k ? 1 + 4 + 1 = 3 ? 5 + 4 + 1 = 9 ? 8 + 4 + 1 = 13 ? 15 + 4 + 1 = 19 ? 29 + 4 + 1 = 36